### v1

# Hierarchical Long Video Understanding with Audiovisual Entity Cohesion and Agentic Search

Abstract
Long video understanding presents significant challenges for vision-language models due to extremely long context windows. Existing solutions relying on naive chunking strategies with retrieval-augmented generation, typically suffer from information fragmentation and a loss of global coherence. We present HAVEN, a unified framework for long‑video understanding that enables coherent and comprehensive reasoning by integrating audiovisual entity cohesion and hierarchical video indexing with agentic search. First, we preserve semantic consistency by integrating entity-level representations across visual and auditory streams, while organizing content into a structured hierarchy spanning global summary, scene, segment, and entity levels. Then we employ an agentic search mechanism to enable dynamic retrieval and reasoning across these layers, facilitating coherent narrative reconstruction and fine-grained entity tracking. Extensive experiments demonstrate that our method achieves good temporal coherence, entity consistency, and retrieval efficiency, establishing a new state-of-the-art with an overall accuracy of 84.1% on LVBench. Notably, it achieves outstanding performance in the challenging reasoning category, reaching 80.1%. These results highlight the effectiveness of structured, multimodal reasoning for comprehensive and context-consistent understanding of long-form videos.
1 Introduction
The rapid growth of long-form video content in domains such as entertainment, education, and surveillance poses significant challenges for automated understanding systems. Unlike short clips, hour-long videos demand reasoning over tens of thousands of frames and audio streams, where events unfold across extended durations and depend on subtle cross-scene relationships. The interleaving of multiple entities, dynamic interactions, and evolving multimodal signals within such content makes coherent interpretation and entity tracking particularly difficult.
Figure 1: The proposed hierarchical video indexing with audiovisual entity cohesion and agentic search.
Recent advances in large Vision-Language Models (VLMs) have improved tasks like captioning, question answering, and temporal grounding on short videos, yet their application to long videos remains constrained by limited context windows and computational bottlenecks. To mitigate these issues, techniques such as adaptive sampling [9, 30, 12], and token compression [24, 22, 21, 15] are proposed, enabling longer sequences under resource limits. Memory-based approaches [5, 8] further extend temporal coverage by dynamically retaining and updating salient information. While these techniques improve scalability, they often sacrifice critical details or contexts, and struggle to preserve the semantic continuity across distant segments.
To address semantic continuity, two new paradigms have emerged: retrieval-augmented generation (RAG) and video agents. RAG-based approaches segment videos into retrievable units and dynamically fetch relevant content during inference, enabling scalable and context-aware comprehension [19, 13, 10, 29, 20]. In parallel, agent frameworks treat large language models as autonomous agents capable of planning, querying, and reasoning over video content [25, 4, 31]. Despite recent progress, RAG-based and agent-based paradigms still face fundamental limitations when applied to long-form video understanding. First, retrieval often depends on local cues such as clip-level captions, producing fragmented and redundant information that undermines semantic coherence. Second, the absence of a hierarchical representation of video content prevents structured reasoning for agents. Finally, without global context, models can only resort to multiple rounds of retrieval to reconstruct continuity, introducing unnecessary complexity.
In this paper, we overcome these limitations by proposing a unified, agentic framework for long video understanding that transitions from fragmented processing to coherent comprehension (see Fig. 1). Our approach is based on two core innovations: audiovisual entity cohesion and hierarchical indexing. First, we introduce a mechanism for entity-level cohesion across visual and auditory modalities, enabling consistent tracking and integration of entities throughout the video. This mitigates semantic fragmentation and preserves narrative continuity. Second, we construct a hierarchical information database that organizes video content at multiple levels—global summary, scene, segment, and entity—allowing for multi-granularity retrieval. Built atop this index, our agentic search mechanism dynamically navigates the hierarchy, enabling goal-driven reasoning. By integrating these components, our framework achieves a holistic understanding of long videos, effectively bridging fragmented information, preserving semantic consistency, and enabling scalable, structured reasoning.
Our contributions are summarized as follows:
  • Audiovisual entity cohesion: we propose a cross-modal entity integration mechanism that maintains semantic consistency across time and modality, improving entity tracking and narrative coherence.
  • Hierarchical indexing with agentic search: we design a multi-level indexing database and apply an agentic search strategy that dynamically queries and reasons over this hierarchy, enabling effective information access.
  • We validate our framework on long video understanding benchmarks, showing superior performance compared to existing baselines. Notably, our method achieves an accuracy of 81.0% on the LVBench benchmark.
2 Related Work
Long Video Understanding with large VLMs
Large Vision-Language Models have been extended to long-form video tasks using techniques such as adaptive sampling [9, 30, 12] and token compression [24, 22, 21, 15], which reduce token overhead by selecting salient frames or merging redundant information across time and modality. These strategies enable longer sequence processing under memory and compute constraints but often sacrifice critical details or incur additional online computation. Memory-based approaches [5, 8] dynamically retain and update salient content, extending the temporal receptive field and supporting reasoning over longer durations. In addition, other approaches such as LongVLM [27] and VideoStreaming [18] leverage hierarchical token aggregation and memory propagation with global-local semantics to support scalable long video understanding. Despite these advances, existing solutions struggle to balance global coherence with local detail, making it difficult to reason across distant scenes or maintain entity continuity over time.
RAG-based Long Video Understanding
RAG has emerged as a promising paradigm for scaling long video understanding by segmenting videos into retrievable units and dynamically fetching relevant context during inference. Recent works have advanced this paradigm in several directions. VideoRAG [10] retrieves aligned video-text chunks to improve factual grounding. VideoRAG [19] explores graph-based RAG over video captions to enhance entity-level semantics. AdaVideoRAG [29] introduces omni-contextual adaptive retrieval, enabling flexible selection of relevant segments through query intent classification. Vgent [20] further develops video graphs that capture temporal dependencies across clips and support structured reasoning over retrieved content. Although RAG-based methods improve retrieval precision over LVLMs, their reliance on fragmented segments and lack of global context limit temporal coherence, thus making complex reasoning tasks challenging.
Long Video Agents
Unlike static RAG-based retrieval, recent work leverages LLMs as autonomous agents for iterative planning, retrieval, and reasoning over long videos, enabling dynamic interaction through tool use and structured search. Representative works include VideoAgent [25], which treats LLMs as agents that iteratively retrieve and interpret video segments; VideoTree [26], which introduces a tree-based representation for adaptive exploration; LVAgent [4], which coordinates multiple agents in multi-round collaboration; and Deep Video Discovery (DVD) [31], which supports tool-augmented search over global and local video content. Additionally, DrVideo [14] reformulates long videos into document-like structures for iterative retrieval-augmented inference. While these frameworks improve interactive reasoning and scalability, their underlying databases remain simple (e.g., frames, clip captions, visual entities) and often require heavily iterative processes to find the answer, leading to high computational cost.
Hierarchical Video Representation
Long video understanding often requires modeling local-global context and hierarchical relationships across scenes, segments, and entities. Prior works address this via local-global aggregation [27, 18], tree-based structure [26], and graph-based retrieval [29, 19], while [31] tracks subjects in a global registry. Although these methods improve semantic coherence, some models incur high online computation and lack a unified offline hierarchical index spanning video, scene, segment, and entity levels—a core feature of our framework for efficient and coherent long video reasoning.
3 Method
3.1 Overview
Recent RAG-based and agent-based frameworks for long-video understanding typically construct their databases using clip-level captions [31] and global entity sets that link disjoint clips [19, 29, 20]. However, this combination of local-clip and global-entity structure does not adequately reflect the inherent hierarchical nature of video content, where scenes, events and entities appear and evolve in alternating patterns. For global or scene-level queries, e.g. “what is the video about?” or “What is the name of the song sung by the third contestant?”, as well as locally ambiguous queries like “what does the protagonist do at 12:00-12:30?", models require long-range contextual information. Relying solely on online retrieval of local fragments can result in incomplete answers or overwhelm the model with redundant and incoherent information. To address this, we propose an offline hierarchical video indexing approach that parses the video into structured representation at the global, scene, segment and entity levels prior to querying. Furthermore, most existing methods focus primarily on the visual modality and may use subtitles as supplementary text, while overlooking the role of speaker identity. Unlike visual appearance or behavior, which may vary across clips, speaker voice remains relatively consistent, offering a robust cue for maintaining long-range identity coherence.
To overcome these limitations, we introduce a framework for hierarchical video indexing and agentic retrieval, which integrates audiovisual entity cohesion and multi-level temporal abstraction, enabling coherent and efficient long video understanding. We aim to perform an offline parsing to construct a structured database with four-level hierarchy:
$\mathcal{D}=\{\tilde{\mathcal{C}},\tilde{\mathcal{E}},\tilde{\mathcal{S}},\tilde{\mathcal{G}}\}$, (1)
where $\tilde{\mathcal{C}}$, $\tilde{\mathcal{E}}$, $\tilde{\mathcal{S}}$, and $\tilde{\mathcal{G}}$ denote the audiovisual segment-level information, canonical audiovisual entities, scene summaries, and global summary, respectively, as shown in Figure 1. Specifically, speaker diarization is leveraged to maintain consistent speaker identities in audio modality across the entire video. Such speaker identities are then fused with visual modality in a local-to-global manner, forming coherent audiovisual entities that align speech and visual appearance across the entire video. Meanwhile, a LLM-based temporal summarization pipeline merges consecutive semantically related segments into scene-level and global abstractions, providing higher-level contextual anchors for global reasoning.
During inference, we employ an agentic reasoning and retrieval approach that performs query-dependent adaptive search over the video indexing $\mathcal{D}$ through a reasoning-driven loop: $\textit{Think}\rightarrow\textit{Act}\rightarrow\textit{Observe}$ (see Figure 1), following the paradigm introduced in [31]. We introduce a suite of multi-level retrieval tools for querying $\mathcal{D}$. Depending on the granularity and complexity of each query, the planner dynamically selects appropriate tools, ranging from global scene browsing, entity querying to local fine-grained search and inspection, to gather the necessary information for accurate response generation.
Figure 2: Overview of our framework. Left: hierarchical database construction. Right: agentic reasoning. The reasoning LLM calls tools iteratively to collect information and answer the question.
3.2 Database Construction
3.2.1 Audio Information Extraction
Audio serves as a crucial carrier of semantic information, conveying narrative flow, speaker identity, and contextual cues that often extend beyond the visual modality. While most existing works merely utilize transcripts or subtitles as textual supplements, they often overlook the long-range consistency provided by speaker identity. Unlike visual appearance or actions, which may vary significally across scenes, speaker voice remains relatively stable, offering a reliable signal for maintaining temporal coherence, especially in dialog-driven content such as documentaries, TV shows, and vlogs.
To leverage this, we employ automatic speech recognition (ASR) and speaker diarization using WhisperX [3], which jointly produce accurate transcripts with timestamps and consistent speaker labels. These annotations enable us to understand not only “what was said and when”, but also “who said it”, a key factor for maintaining entity-level coherence across time.
3.2.2 Segment Information Extraction
We uniformly divide the entire video into fixed-length temporal segments, which serve as the basic unit of local information. For each video segment $i$, we first extract the corresponding audio content, including speaker labels $P_{i}$ and timestamped transcripts $T_{i}$. These audio features, along with the video frames of the segment, are fed into a VLM to generate a segment-level caption $V_{i}$ and fused speaker information $P_{i}^{\prime}$ that associates visual cues with speaker labels (e.g., appearance, actions, identity for a speaker label). We then construct the textual representation of the segment $C_{i}^{t}$ by concatenating the fused speaker information $P_{i}^{\prime}$, transcript $T_{i}$, and visual caption $V_{i}$:
$C_{i}^{t}=[P_{i}^{\prime}\,;\,T_{i}\,;\,V_{i}].$ (2)
This representation serves as the foundation for subsequent entity-level cohesion and cross-temporal reasoning.
As captioning alone may omit critical visual details, we further incorporate a visual embedding $C_{i}^{v}$ using the universal multimodal retrieval method UNITE [11] to enrich the representation. The final unified segment-level representation is defined as
$C_{i}=[C_{i}^{t}\,;\,C_{i}^{v}],$ (3)
which grounds each utterance and provides localized visual context for each speaker, composing the segment database $\tilde{\mathcal{C}}$. In our implementation, each segment spans 30 seconds, and we sample 20 frames per segment for caption generation.
3.2.3 Audiovisual Entity Extraction
For each segment $i$, we employ a LLM to extract key entities from its textural representation $C_{i}^{t}$. The LLM identifies and categorizes salient subjects like characters, locations, and events, producing a set of entities with summarized descriptions:
$E_{i}=\{e_{1}^{i},e_{2}^{i},\ldots,e_{N_{i}}^{i}\},$ (4)
where each $e_{k}^{i}$ represents a semantic entity, including its name and description derived from the audiovisual context of segment $i$.
While prior approaches often merge entities using simple name matching or cosine similarity thresholds [19, 20], such heuristics are insufficient for long videos with complex variations. For example, a character appearing in different scenes with varying outfits (e.g. red dress vs. yellow dress) or emotional states (e.g. happy old man vs. angry male teacher) may be incorrectly split into multiple entities, while two visually similar but distinct characters may be mistakenly merged.
To address this, we perform entity consolidation in two stages, embedding-based clustering followed by LLM-based canonicalization. First, all entity descriptions $\{e_{k}^{i}\}$ are embedded using a text encoder $f_{\text{text}}(\cdot)$ to obtain dense representations: $z_{k}^{i}=f_{\text{text}}(e_{k}^{i}).$ We then apply clustering to group semantically similar entities into preliminary clusters, capturing potential cross-segment correspondences. Second, each cluster is re-evaluated by an LLM, which reads all member entities and either generates a unified canonical summary or separates them into distinct subgroups if semantic conflicts are detected. This LLM-in-the-loop verification prevents both over-merging and over-splitting, resulting in a refined set of canonical entities:
$\tilde{\mathcal{E}}_{g}=\{\tilde{e}_{1},\tilde{e}_{2},\ldots,\tilde{e}_{J}\},$ (5)
where each $\tilde{e}_{j}$ aggregates temporally distributed mentions of the same underlying identity or concept.
To further enhance coherence, we emphasize speaker label cues into the consolidation process. If multiple segments share the same speaker label, their corresponding character entities about characters are prioritized for merging, even when visual or textual features differ. This enables robust audiovisual entity cohesion, ensuring consistent identity tracking across the video. For instance, a host narrating in both indoor and outdoor scenes will be recognized as the same entity based on voice consistency, rather than being fragmented due to appearance changes.
After consolidation, each canonical entity $\tilde{e}_{j}$ is associated with a global description and a set of linked segments $Q_{j}$. During retrieval, incorporating all linked segments of top-$K_{1}$ entities can be computationally expensive. Moreover, these segments often contain extraneous query-irrelevant details, degrading cosine similarity matching between the query and segment embeddings, while LLM-base reranking incurs high computational cost. To address this, we further introduce an entity-centric re-captioning process on original segment textual representation during offline database construction. For each linked segment $i$ and entity $\tilde{e}_{j}$, we generate a focused description $\tilde{C}^{t}_{i,j}$ using an LLM. This description centers on the entity’s appearance, actions, and events within the segment, while excluding irrelevant context. The final entity database consists of $\tilde{\mathcal{E}}_{g}$ and this fine-grained entity-segment links $\tilde{\mathcal{E}}_{e}=\{\tilde{C}^{t}_{i,j}|j=1,2,\ldots,J,i\in Q_{j}\}$, i.e. $\tilde{\mathcal{E}}=\{\tilde{\mathcal{E}}_{g}\,;\,\tilde{\mathcal{E}}_{e}\}$. During retrieval, after initial entity matching in embedding space of $\tilde{\mathcal{E}}_{g}$, we re-rank candidate segments by computing similarity between the query and $\tilde{C}^{t}_{i,j}$, selecting the top-$K_{2}$ segments for precise entity grounding. This approach balances retrieval accuracy and computational efficiency, avoiding redundant noise and excessive LLM overhead.
3.2.4 Scene Segmentation and Global Summary
Long videos often consist of multiple temporally extended scenes that share coherent narratives, recurring characters, and consistent environments. Existing approaches typically segment videos into very short clips (e.g., 5 seconds in [31]) or aggregate fixed-size chunks [19], which may capture isolated moments but fail to model long-range temporal dependencies. To address this, we perform adaptive scene-level aggregation based on the segment-level descriptions $\{C_{i}^{t}\}_{i=1}^{N}$, where $N$ is the total number of segments. These captions are split into overlapping chunks and processed by an LLM to perform semantic grouping, merging consecutive segments into coherent scenes:
$\mathcal{S}=\{s_{j}\}_{j=1}^{M},\quad s_{j}=\{c_{a_{j}},c_{a_{j}+1},\ldots,c_{b_{j}}\},$ (6)
where each $s_{j}$ represents a temporally contiguous scene with a consistent narrative focus, and $c_{i}$ denotes the $i$-th segment. The boundaries $(a_{j},b_{j})$ are adaptively determined by the LLM based on the continuity of events observed across captions.
Each resulting scene $s_{j}$ is then summarized by the LLM to produce a concise scene-level description $\tilde{s}_{j}$, capturing its key characters, events, and transitions. The collection of these summaries forms the high-level scene set $\tilde{\mathcal{S}}=\{\tilde{s}_{1},\tilde{s}_{2},\ldots,\tilde{s}_{M}\}$, which provides a structured narrative view of the video. Finally, a global summary $\tilde{\mathcal{G}}$ is generated from $\tilde{\mathcal{S}}$, describing the main storyline, recurring entities, and overall context (e.g., background and video type).
3.3 Agentic Search with Reasoning
Similar to [31], our agentic search framework adopts an iterative think–act–observe loop for autonomous and adaptive information retrieval. At the core, we use a reasoning LLM as the planner. The reasoning process is tool-augmented: the agent first plans actions and formulates queries (think), then invokes specialized tools with appropriate parameters (act), and finally integrates tool outputs into subsequent reasoning (observe). This synergy between LLM-driven reasoning and tool interaction enables progressive query refinement, evidence-guided navigation, and accurate answer generation.
To provide holistic context, the global summary $\tilde{\mathcal{G}}$ is prepended to each initial prompt, while tool outputs dynamically enrich the agent’s reasoning across iterations. This design highlights the critical role of tools in extending the agent’s capabilities and underscores the importance of tight LLM–tool collaboration for robust and adaptive retrieval.
3.3.1 Multi-Granularity Tools
We denote the set of all available tools as
$\mathcal{T}=\{T_{\text{scene}},T_{\text{caption}},T_{\text{visual}},T_{\text{entity}},T_{\text{inspect}}\}.$ (7)
Each tool $T_{i}\in\mathcal{T}$ is a callable function that maps a textual query (or question) and optional time ranges to a textual response and corresponding timestamps $T_{i}(q,\tau;D)\rightarrow(r,\tau^{\prime})$, where $q$ denotes the user query, $\tau=[t_{s},t_{e}]$ the optional input time window, $D$ the corresponding database, $r$ the retrieved or generated response, and $\tau^{\prime}$ the associated timestamps.
Global Scene Browse.
The Global Scene Browse tool $T_{\text{scene}}$ supports coarse-grained navigation and scene localization along the video timeline. Given a user query $q$ and the scene collection $D=\tilde{\mathcal{S}}$, it identifies and summarizes the most relevant scenes with an LLM, returning their storyline and corresponding timestamps $\tau^{\prime}$. The agent tends to invoke this tool for complex or ambiguous queries involving multiple events or temporal dependencies.
Segment Caption Search.
The Segment Caption Search tool $T_{\text{caption}}$ performs fine-grained text-based retrieval within specified temporal ranges. Given the user query $q$ and the segment database $D=\tilde{\mathcal{C}}$, the tool retrieves the most semantically relevant segment descriptions $r$ along with their associated time spans $\tau^{\prime}$. This is achieved through cosine similarity matching between the query embedding and pre-computed caption embeddings for all video segments, ensuring efficient and accurate retrieval of localized content.
Segment Visual Search.
To capture visual cues that may be overlooked in textual descriptions, the Segment Visual Search tool $T_{\text{visual}}$ serves as a complementary mechanism to Segment Caption Search. While the latter relies on text embeddings, $T_{\text{visual}}$ leverages cross-modal embeddings generated by the UNITE framework [11], as previously described. This enables retrieval based on rich visual semantics aligned with the query, ensuring that visually salient details are incorporated into the search process.
Entity Search.
The Entity Search tool $T_{\text{entity}}$ supports high-level, entity-centric retrieval across large temporal ranges. Given an entity-related query $q$, the tool first retrieves the top-$K_{1}$ most relevant entities from database $\tilde{\mathcal{E}_{g}}$ based on their descriptions in the pre-computed embedding space. For all segments linked to these entities, it then performs a second-stage reranking to select the top-$K_{2}$ most relevant segments from the entity-centric database $\tilde{\mathcal{E}_{e}}$, using the same query. Finally, the tool applies entity-aware re-captioning to the retrieved segments, generating a coherent response $r$ enriched with precise timestamps $\tau^{\prime}$.
Inspection Tool.
The Inspection tool $T_{\text{inspect}}$ provides fine-grained temporal inspection to support detailed reasoning. It consists of two complementary modules: Clip Caption Inspect ($T_{\text{inspect}}^{\text{tex}}$) and Visual Inspect ($T_{\text{inspect}}^{\text{vis}}$). The $T_{\text{inspect}}^{\text{tex}}$ examines coarse textual descriptions to determine what occurred during a specified time span. For example, for a query such as “What does the protagonist do after he jumps down the stairs?”, the tool inspects subsequent time ranges to identify the protagonist’s next actions after locating the event “he jumps down the stairs” in previous iteration. The $T_{\text{inspect}}^{\text{vis}}$ tool, leverages a VLM to perform precise visual verification within a given time range. Due to frame limitations in the VLM we used, this inspection focuses on short intervals, ensuring accurate visual grounding for fine-grained queries.
In both segment search and inspection tools, we prioritize low-cost text-based retrieval over high-cost visual reasoning with VLMs, enabling our system to achieve efficient and coherent long-term video understanding. Together, these multi-granularity tools effectively support diverse query types across our hierarchical database, enabling flexible and accurate information access.
3.3.2 Multi-Step Reasoning
Besides $(\mathcal{T},D)$, the agent maintains a context memory $\mathcal{M}$ for the reasoning process, initiated to include the global summary $\tilde{\mathcal{G}}$, i.e. $\mathcal{M}_{0}=\{\tilde{\mathcal{G}}\}$. Let $q_{org}$ denotes the original question. At each reasoning step $t$, the planner dynamically selects to call a tool $T_{a_{t}}$ with a modified query $q_{t}$ with optional time ranges based on the current context: $a_{t}=\pi_{\text{LLM}}(q_{org},\mathcal{T},\mathcal{M}_{t})$, where $\pi_{\text{LLM}}$ denotes the LLM-driven policy that determines which tool to invoke. The tool results $(r_{t},\tau^{\prime}_{t})=T_{a_{t}}(q_{t};D)$ and the tool call context are integrated into the context memory to guide subsequent reasoning within the think–act–observe loop, i.e. $\mathcal{M}_{t}=\{\mathcal{M}_{t-1};(T_{a_{t}},q_{t});(r_{t},\tau^{\prime}_{t})\}$. The iteration continues until it finds the answer or the maximum iteration step is reached. Figure 2 shows an example for a counting-like query that needs long contexts. The agent first calls the Global Scene Browse tool to locate the “second-to-last song” and obtain its content and time span. It then calls Visual Inspect tool to analyze the segment of the song “One" (55:00–57:00), which returns “two singers playing guitar and one playing percussion", enabling the agent to derive the final answer.
4 Experimental Results
Table 1: Comparison results on LVBench. Accuray (%) is reported. Bold indicates the best performance, and underscore denotes the second-best.
Methods ER EU KIR TG Rea Sum Overall
Proprietary VLMs
GPT-4o [1] 48.9 49.5 48.1 40.9 50.3 50.0 48.9
OpenAI o3 [17] 57.6 56.4 62.9 46.8 50.8 67.2 57.1
Open-source VLMs
Qwen2.5-VL-72B [2] - - - - - - 47.7
AdaReTake [24] 53.0 50.7 62.2 45.5 54.7 37.9 53.3
Seed1.5-VL-Thinking-200B [7] 65.4 63.4 68.0 53.6 63.7 46.6 64.6
RAG and agents
VideoRAG [19] 47.4 49.3 57.1 36.5 43.9 39.7 49.2
VideoTree [26] 30.3 25.1 26.5 27.7 31.9 25.5 28.8
VideoAgent [25] 28.0 30.3 28.0 29.3 28.0 36.4 29.3
VideoLucy [32] 54.3 59.8 75.6 51.7 55.9 49.1 58.8
DVD [31] 73.4 73.3 80.4 72.3 70.7 74.1 74.2
DVD w. subtitle [31] 75.5 77.1 79.0 72.7 68.7 84.5 76.0
Ours 79.9 82.6 84.0 86.7 79.6 79.3 81.0
Ours (2 fps) 83.2 84.8 88.2 82.0 80.1 84.5 84.1
Table 2: Comparison results on various long video benchmarks. We use 2 fps in captioning for LVBench and 0.67 fps for other benchmarks.
Methods LVBench Video-MME (long) LongVideoBench EgoSchema
Overall w/o subs w subs Long-val Val Val
Proprietary VLMs
GPT-4o 48.9 65.3 72.1 60.9 66.7 70.4
OpenAI o3 57.1 64.7 - 60.6 67.5 63.2
Open-source VLMs
Qwen2.5-VL-72B 47.7 63.9 - - 60.7 -
AdaReTake 53.3 65.0 76.4 - 67.0 -
Seed1.5-VL-Thinking-200B 64.6 - - - 74.0 -
RAG and agents
VideoRAG 49.2 - - - - -
VideoTree 28.8 - - - - 67.0
VideoAgent 29.3 - - - - 63.2
VideoLucy 58.8 66.8 - - - -
DVD 74.2 67.3 - 68.6 71.6 76.6
DVD w. subtitle 76.0 - - - - -
Ours 84.1 - 85.5 81.0 83.0 81.6
4.1 Benchmark and Implementation Details
Benchmark
We evaluate our approach on four widely-used long video understanding benchmarks, including LVBench [23], Video-MME [6], LongVideoBench [28] and EgoSchema [16]. The LVBench, a standard benchmark for long-term video question answering, contains 1,549 questions across 103 videos, with an average duration of 4,101 seconds. The dataset covers six diverse categories: temporal grounding (TG), summarization (Sum), reasoning (Rea), entity recognition (ER), event understanding (EU), and key information retrieval (KIR). With high-quality ground-truth annotations, LVBench provides a reliable and challenging benchmark for evaluating comprehensive long video understanding capabilities. For Video-MME, we evaluate it on its long split, which includes 300 videos and 900 questions, with durations ranging from 30 to 60 minutes. For LongVideoBench, we focus on the long subset in the val split, containing 188 videos and 564 questions, with durations between 900 and 3600 seconds. Most videos in this subset lack audio tracks. It is noted that we utilize the official subtitles without speaker identity on LongVideoBench. EgoSchema serves as a diagnostic benchmark for long-form understanding and reasoning. We evaluate on its val split, which contains 500 videos and 500 questions, with each video lasting three minutes and no audio provided. For simplicity, we use audio streams only when the language is English across all datasets.
Implementation Details
For database construction, we employ GPT-4.1 [1] to generate segment-level captions and summarize scenes and entities. Each 30-second segment is captioned using 20 sampled frames (0.67 fps) to ensure efficiency while maintaining coverage. During agentic search, we use OpenAI o3 [17] as the reasoning planner, which aggregates information from tool outputs and produces answers, with a maximum reasoning depth of 10 steps. For query-aware re-captioning during tool calls, we sample 30 frames per segment (1 fps) to provide sufficient visual context. Additionally, we leverage OpenAI o3 as the VLM used in Visual Inspect tool and set its maximum number of input frames to 50. Unless otherwise specified, the same settings are applied across all datasets.
Table 3: Ablation study of various modules on LVBench. All variants use 0.67 fps in captioning.
Methods Speaker Transcript Hierarchical Visual embed Acc.
Ours_clip_t 72.1
Ours_clip 72.8
Ours_visual 71.7
Ours_transcript 75.7
Ours 81.0
4.2 Comparison with State-of-the-Art
We compare our framework against several leading approaches for long video understanding, including proprietary large VLMs [1, 17], open-source VLMs [2, 24, 7], RAG-based systems [19] and video agents [26, 25, 32, 31]. We reproduce the results of VideoRAG [19] on LVBench using the official implementation, while other baseline results are taken directly from published reports.
Table 1 summarizes the comparison results on LVBench. Our approach consistently outperforms all baselines in overall accuracy, with particularly strong gains in reasoning (Rea, a challenging category in all baselines) and temporal grounding (TG). These improvements highlight the effectiveness of our multi-granularity information parsing during offline database construction and the multi-level retrieval tools integrated into our agentic search framework. In entity recognition (ER), our method also surpasses DVD, which relies on global subject registries, demonstrating the robustness of our audiovisual entity cohesion strategy. Notably, our system achieves these results using much coarser temporal segmentation (30-second segments with 20 frames for captioning) and fewer reasoning iterations (maximum of 10) compared to DVD’s finer-grained configuration (5-second segments at 2 fps) and deeper reasoning depth (up to 15 steps), underscoring both the efficiency and scalability of our approach.
As our approach employs coarse temporal segmentation and sampling than DVD, our captions inherently contain fewer details. To assess the impact of this difference in textual richness, we compare our scheme with a variant that samples at 2 fps for each 30‑second segment (denoted as Ours (2 fps) in Table 1). Specifically, this variant generates captions for two 15-second sub-segments (each sampled at 30 frames) and concatenates them into a single segment description. We can see that with a larger sampling rate, the proposed scheme performs consistently better in all categories, owing to more detailed textual descriptions by captioning. We find that increasing the sampling rate consistently improves performance across most categories, attributable to the richer textual descriptions generated by denser captioning.
Table 2 further presents the comparison results on all benchmarks. Our method achieves the highest scores on all benchmarks, reaching 84.1 on LVBench, 85.5 on Video-MME (long), 81.0/83.0 on long/val split of LongVideoBench, and 81.6 on EgoSchema, even though most videos in LongVideoBench and EgoSchema lack audio streams. These results underscore the effectiveness of reasoning across the proposed hierarchical structure with semantic coherence.
Figure 3: Comparison of accuracy and efficiency on six categories. “Acc" and “iter" denote accuracy and the average number of reasoning iterations, respectively.
Figure 4: Case study on different reasoning chains with tool calls.
4.3 Ablation Study
4.3.1 Ablation on Various Modules
We conduct an ablation study on LVBench to evaluate the contributions of several key components: Audio (Speaker and Transcript), Hierarchical, and Visual embed. “Speaker" and “Transcript" denote the use of speaker labels and transcripts in audio, respectively. “Hierarchical" represents the hierarchical organization of video semantics (segment, entity, scene, and global summary) and “Visual embed" corresponds to the visual retrieval capability enabled by the Segment Visual Search tool.
We compare four variants in Table 3. Ours_clip retains only the segment-level search and two inspection tools, emphasizing reasoning via local grounding. Ours_clip_t further disables the Segment Visual Search from Ours_clip, focusing exclusively on textual retrieval at segment level. Ours_visual removes audio transcripts and speaker annotations while keeping all other components, allowing us to assess the effectiveness of purely visual–structural reasoning. Ours_transcript removes speaker identities by the diarization step and constructs the hierarchical database without any speaker-related information.
As shown in Table 3, removing the hierarchical organization in Ours_clip leads to a substantial performance drop compared to Ours, with accuracy dropping from $81.0$ to $72.8$. This underscores the central role of hierarchical indexing in aggregating multi-granular evidence and supporting effective cross‑segment reasoning. The audio modality also contributes significantly to overall performance. In Ours_visual, the accuracy drops to $71.7$ due to the loss of transcripts and speaker information that not only provide complementary context beyond visual cues but also enhance entity consistency. Indeed, transcripts alone provide meaningful semantic enrichment: Ours_transcript improves by 4% over Ours_visual, indicating that textual audio cues facilitate richer reasoning. Furthermore, incorporating speaker identity yields an additional 5.3% accuracy gain when comparing Ours with Ours_transcript. These results emphasize the effectiveness of both transcripts and speaker labels. Although most questions in LVBench focus on visual content, the audio stream nonetheless strengthens entity coherence, thereby improving semantic consistency within the hierarchical structure. Lastly, Ours_clip_t performs slightly worse than Ours_clip, because it lacks visual embedding-based search, which offers additional grounding capabilities. All these results demonstrate that these components contribute uniquely to the system’s overall effectiveness.
4.3.2 Analysis on Six Categories
To further assess the effectiveness of our method across different query types, we compare Ours with Ours_clip in terms of both accuracy and the average number of reasoning iterations required to answer queries across all six categories. As shown in Figure 3, our approach consistently achieves higher accuracy with fewer iterations compared to the non-hierarchical baseline, with particularly notable gains in entity recognition and reasoning. This improvement is expected, as hierarchical indexing provides readily accessible information at multiple granularities, enabling the agent to retrieve relevant context more efficiently and thereby reducing reasoning steps. It is worth noting that the summarization category in LVBench often involves fine-grained, step-level queries (e.g., “What does Mandy do after she stands in front of the judges’ table?”), which inherently require multi-turn reasoning rather than simple scene summarization.
4.4 Qualitative Analysis
We present several qualitative examples in Figure 4 to illustrate the agent’s adaptive reasoning behavior across diverse question types. For global queries requiring holistic understanding (e.g., the athletes’ nationality), the agent retrieves the answer directly from the global summary without invoking additional tools. In contrast, for queries involving specific entities (e.g., the vending machine or the boy), the Entity Search tool is first employed to localize the relevant visual context, followed by targeted inspection to infer detailed attributes or actions. For queries demanding fine-grained visual cues (e.g., numerical content on a webpage) that may not be captured in captions, the Segment Visual Search tool effectively supplements missing information. Finally, when the query specifies an explicit temporal scope and sufficient contextual evidence is available, the Visual Inspect tool is utilized to verify and refine the final answer. These examples highlight the model’s capability to dynamically compose tool chains according to question type and evidence distribution across modalities, demonstrating robust adaptability in complex reasoning scenarios.
5 Conclusion
In this work, we tackled the challenges of long video understanding by proposing a unified framework that integrates offline hierarchical video indexing with agentic multi-granularity retrieval. Our approach organizes video content across four levels, global, scene, segment, and entity, while incorporating audiovisual entity cohesion to maintain semantic consistency over extended temporal spans. Extensive experiments on LVBench demonstrate that our method significantly outperforms state-of-the-art baselines in overall accuracy, highlighting the effectiveness of hierarchical indexing and multimodal entity integration. These findings underscore the potential of structured, multi-level representations for advancing long video comprehension, a promising direction for future research.
Appendix A Visualization of Entity Coherence by Auditory Clues
In this section, we further illustrate the role of speaker identity in entity tracking. Figure 5 presents two examples. In these cases, characters undergo dramatic appearance changes, such as clothing, viewpoint shifts, and occlusion, throughout the video, complicating the merging of fragmented descriptions into a single entity. Leveraging speaker coherence, however, enables successful consolidation of these descriptions.
Figure 5: Cases: speaker identity for entity tracking.
Appendix B Additional Case Study
As shown in Figure 6, we present more complex question-answer cases that involve multiple reasoning steps and tool calls, further demonstrating the effectiveness of our proposed pipeline. For the interaction question involving the appearance of the book “Zurked” with a thumbs-down, the agent first performs a coarse segment-level caption search, then progressively narrows the temporal window via entity search. After validating the candidate clips with caption inspection in a long time range, it applies fine-grained visual inspection to extract the correct interaction outcome. For the reasoning-oriented question “Why does Brynn Cummings’ mother wipe away tears?”, the agent again initiates with segment caption search tool, but due to incomplete textual matches, it escalates to segment-level visual search to gather more reliable evidence. Once the event is localized, the agent conducts a targeted visual inspection to infer the emotional cause behind the mother’s reaction.
Figure 6: Cases with multiple reasoning steps.
Appendix C Prompt for the Planner
We present the planner’s prompt for agentic search in Table 4. This prompt guides the planner in selecting the most appropriate tools to search from the hierarchical database and determining what information to request at each reasoning step, thereby enabling systematic information gathering and progressively moving toward the final answer.
Table 4: The agentic search prompt structure. The prompt is divided into four distinct sections: goal, tools, tool preferences and hints.
Module Prompt Content
GOAL You will be given a set of tools to assist you exploring the video, understanding it and reasoning the answer.
Please follow the THINK $\rightarrow$ ACT $\rightarrow$ OBSERVE loop:
$\bullet$ THOUGHT: Reason step-by-step about what question to ask and which tool to call next.
$\bullet$ ACTION: Call exactly one tool that moves you closer to the final answer.
$\bullet$ OBSERVATION: Summarize the tool call’s output.
Continue the loop until the user’s query is fully resolved, then end your turn with the final answer.
TOOLS Here are tools you can use to reveal your reasoning process whenever the provided information is insufficient:
$\bullet$ global_scene_browse_tool: for scene-related query to explore scenarios, temporal orders, and contextual structure in a rough manner without precise details (e.g., first appearance, second song, third collision).
$\bullet$ entity_search_tool: for entity-related information retrieval, finding important subjects involved in events.
$\bullet$ clip_caption_search_wtime_tool: to search from rough captions and audio transcriptions of local clips within a list of time ranges related to a query. If you want to search from the whole video, use the whole video time range.
$\bullet$ clip_visual_search_wtime_tool: to search from visual features of local clips within a list of time ranges (list[tuple[HH:MM:SS, HH:MM:SS]]) related to a query. If you want to search from the whole video, use the whole video time range. This tool may provide more detailed information as a supplement to clip_caption_search_wtime_tool.
$\bullet$ clip_caption_inspect_tool: to extract rough captions and audio transcriptions of local clips within any list of time ranges (list[tuple[HH:MM:SS, HH:MM:SS]]). This tool is suitable for further inspecting what happened in a time range before or after some events happened.
$\bullet$ visual_inspect_tool: to extract details from local visual clips within a narrow list of time ranges that covers less than 50 seconds to answer a question or retrieve query-related details. When the time ranges cover over 50 seconds, first use clip_caption_inspect_tool to get a rough context.
$\bullet$ finish: Once you believe you have found the answer, you can call the finish tool with an answer.
TOOL PREFERENCES $\bullet$ When no context is given, call global_scene_browse_tool for scene-related queries to get an overview of related context and timelines,
or call clip_caption_search_wtime_tool for event-related queries to get a rough context and timelines,
or call entity_search_tool for entity-related queries.
$\bullet$ When you cannot locate the needed context from scene or entity tools, use clip_caption_search_wtime_tool or clip_visual_search_wtime_tool to expand your search.
$\bullet$ If the retrieved material by clip_caption_search_wtime_tool lacks relevant contexts, further call clip_visual_search_wtime_tool for more fine-grained search.
$\bullet$ If the retrieved material lacks precise, question-relevant detail (e.g., an unknown name, count) or you are uncertain of an answer after searching, call clip_caption_inspect_tool
or inspect frames with visual_inspect_tool with a list of time ranges to take a closer look.
$\bullet$ After locating an answer in the script, always make a CONFIRM with visual_inspect_tool query.
HINTS $\bullet$ Before giving the final answer, confirm critical visual or numeric facts with visual_inspect_tool.
$\bullet$ If you call clip_caption_search_wtime_tool in three consecutive times but still cannot find useful information to answer the question, please try clip_visual_search_wtime_tool to get more detailed information.
[…]
$\bullet$ To make a good plan to questions that need complex reasoning, sometimes you need to first ask some other related contexts instead of directly asking the target question,
$\bullet$ For questions that need counting the number of times an event occurs over time, call global_scene_browse_tool first.
If you are uncertain about its answer, please search for related events or subjects without counting first to find all related information, then do the counting based on observations.
$\bullet$ Your final answer must be concise and directly address the question.